OpenStudy (anonymous):

4(2x - 3) + 8(x - 4) = 2(2x + 6) in fraction form

OpenStudy (anonymous):

solve for x

OpenStudy (hoblos):

x=14/3

OpenStudy (lilg132):

how did you get that

OpenStudy (anonymous):

yeah...

OpenStudy (hoblos):

first expand =>8x -12+8x-32 = 4x +12 =>16x - 44 = 4x +12 =>12x=56 =>x=56/12 = 14/3

OpenStudy (lilg132):

ok i see where i made a mistake thanks

OpenStudy (hoblos):

welcome :D

OpenStudy (anonymous):

okay thanks